OpenStudy (haleholmes):

Evaluate: log6 (1/36) Please help me understand.

OpenStudy (haleholmes):

\[\log_{6} (1/36)\]

OpenStudy (freckles):

hint 36=6^2

OpenStudy (freckles):

use the division property for logs

OpenStudy (haleholmes):

I have almost zero understnading you'll have to walk me through it. My math lesson din't really help.

OpenStudy (freckles):

\[\log_a(\frac{n}{m})=\log_a(n)-\log_a(m) \\ \log_a(1)=0 \\ \log_a(a)=1 \\ \log_a(x^r)=r \log_a(x)\]

OpenStudy (freckles):

i would use all of these properties

OpenStudy (freckles):

it is like a puzzle

OpenStudy (freckles):

try to see if you can fit these together somehow to get an answer

OpenStudy (haleholmes):

For evaluating should i get a single number or should I still end up with an equation?

OpenStudy (haleholmes):

I have NO idea

OpenStudy (freckles):

Well you don't have an equation to start out with... You have a expression so you will end up with an expression. It will be a more simplified expression.

OpenStudy (freckles):

do you know how to use the first property I mentioned?

OpenStudy (haleholmes):

No, I need a walk through

OpenStudy (haleholmes):

None of it makes sense

OpenStudy (freckles):

\[\log_a(\frac{n}{m})=\log_a(n)-\log_a(m) \\ \text{ and you have } \\ \log_6(\frac{1}{36})\] so it looks like n is 1 and m is 36 and a is 6 I bet you can plug in those numbers.

OpenStudy (haleholmes):

\[\log_{6}(1) - \log_{6}(36) \]

OpenStudy (freckles):

right

OpenStudy (freckles):

\[\log_a(\frac{n}{m})=\log_a(n)-\log_a(m) \\ \log_a(1)=0 \\ \log_a(a)=1 \\ \log_a(x^r)=r \log_a(x) \] now looking at these equations here which one can be applied to the log(1) thing?

OpenStudy (haleholmes):

The second one. So it is 0 now?

OpenStudy (freckles):

so then we just have to look at \[-\log_6(36)\]

OpenStudy (freckles):

and the hint I gave is that 36 can be written as 6^2

OpenStudy (freckles):

that 2 is a power which equation has a power involved from my list

OpenStudy (haleholmes):

the last one

OpenStudy (freckles):

\[- \log _6(6^2)=-(2)\log_6(6)\] so you see the power can be brought down from the last one

OpenStudy (haleholmes):

Yup.

OpenStudy (freckles):

now you have one more property to use and you are done

OpenStudy (freckles):

look at the list one more time

OpenStudy (haleholmes):

The log6(6)= 1

OpenStudy (haleholmes):

So the answer is -2?

OpenStudy (freckles):

\[\log_6(\frac{1}{36}) \\ \log_6(1)-\log_6(36) \\ 0-\log_6(36) \\ -\log_6(36) \\ -\log_6(6^2) \text{ since } 36=6^2 \\ -(2)\log_6(6) \\ -2(1) \\ -2 \\ \text{ and yes }\]

OpenStudy (anonymous):

yes

OpenStudy (haleholmes):

Thanks so much! That made sense. :D